Hi, I have a question and it is find the slope of the linear function
f(t)=(3-4t)/5.
I'm not sure, but this is what I came up with:
f(t)=(3-4t)/5
5=3-4t
f(t)= -4t-2
Therefore since it's in form of y=mx+c the slope is m, which is -4
Am I even on the right track?????
Thanks, Christa.

Answer by venugopalramana(3286)

SEE MY COMMENTS
Hi, I have a question and it is find the slope of the linear function
f(t)=(3-4t)/5.
I'm not sure, but this is what I came up with:
f(t)=(3-4t)/5.....OK
5=3-4t....NO YOU FORGOT TO TAKE F(T) ON LHS..IT IS
5F(T)=3-4T.....BUT THIS IS NOT THE WAY...HAVE A CLEAR OBJECTIVE IN VIEW..
STD.EQN.IS Y=MX+C...WHERE SLOPE=M AND Y INTERCEPT=C
SO.....CALL F(T)=Y
Y=(3-4T)/5=15-4T/5
HENCE M= SLOPE=-4/5